2.11 Projecting Transit Ridership. As a transit planner your job is to predict ridership and total fare revenue. Suppose the short-run elasticity of demand for com muter rail (over a 1-month period) is 0.60, and the long-run elasticity (over a 2-year period) is 1.60. The current ridership is 100,000 people per day. Suppose the transit authority decides to increase its fares from $2.00 to $2.20. a. Predict the changes in train ridership over a 1-month period (the short run) and a 2-year period (the long run). b. Over the 1-month period, will total fare revenue increase or decrease? What about the 2-year period?
Answer #1

Elasticity of demand = % Change in quantity demanded / % Change in price

% Change in price = [$(2.2 - 2) / $2] x 100 = ($0.2 / $2) x 100 = 10%

(a)

(i) In short run,

0.6 = % Change in quantity demanded / % Change in price

% Change in quantity demanded = 0.6 x 10%

% Change in quantity demanded = 6% (A decrease, following law of demand)

New ridership = 100,000 x 94% = 94,000 (Decrease by (100,000 - 94,000) = 6,000)

(ii) In long run,

1.6 = % Change in quantity demanded / % Change in price

% Change in quantity demanded = 1.6 x 10%

% Change in quantity demanded = 16% (A decrease, following law of demand)

New ridership = 100,000 x 84% = 84,000 (Decrease by (100,000 - 84,000) = 16,000)

(b)

Since elasticity is less than 1 in short run, demand is inelastic and increase in price will increase total revenue.

Since elasticity is higher than 1 in long run, demand is elastic and increase in price will decrease total revenue.